plastic doll - 1. what is the difference detween energy and power? what is a unit of power? How dose speed relate to power?

Explain how a sound is actually a form of energy? in what medium do sound is actually a form of enery ? In what medium do sound waves travel?

3. Give an example of change of enery from potential to kinetic; from kinetic to potental.

4. What is a trophic level ? Give some examples. How much energy is lost at each trophic level?

5. If you eat 600 calories perday [roughly one dessert ] below your energy needs, how long would it take tolose 10 pounds? how long would it take to lose 10 pounds? How long would you have to walk [assuming 80 calories burned per mile walked] to lose 10 pounds?

6. What make a good conductor of heat? What makes a good insulator?

7. Expain why heat only flows spontaneously from a hot to cold object.

8. In what way is aging an example of the second law of thermodynamics?
Answer #1
ENERGY POWER
1.it is capasity to do work. 1.it is rate of doind work
2.S.I. units of energy is JOULS.

2.S.I. units of power is WATT

1 WATT = 1JOULS/SECOND

3.energy is usage of power over a time 3.power is used to operate all devices

4.we cannot generate energy.

energy can be converted from one form to the another form

4.we can generate power

we can generate power

5.we all ways loose some energy in the process.

the amount of energy we retain is called efficiency

4.power is consumed at the same time that is generated

6.energy can be stored.

but electricity is not easy to store.electricity have to convert to another form like chemical, physical and gravitational.

5.we cannot store or save power

2. POWER = FORCE \times SPEED

FORCE = MASS \times ACCELERATION

POWER = MASS \times ACCELERATION \times SPEED

sound is a form of energy made by vibrations

when an object vibrates it causes air particle arround it to move.

sound waves require some material medium to travel through

sound waves can travel through solids liquids and gases

sound waves can not travel through vaccume

3. energy can be changed but it never lost

when we have thrown a ball initially it has maximum kinetic energy and minimum potential energy

when that ball reaches the maximum height it has maximum potential energy and minimum kinetic energy

when that boll reaches the ground it has maximum kinetic energy and minimum potential energy

as a skier skis up the hill kinetic energy is lost and potential energy is gained

as the skier skis down the hill potential energy is lost and the kinetic energy is gained

4. each step in a food chain is called a trophic level

primary producers make up the first trophic level.

example :plants and grass

energy is lost at each trophic level

energy is unavailable for the next trophic level

energy is lost through respiration and then converted to heat, stored in dead organisms and in waste materials

there fore only avalable in decomposers

because of this there is less living tissue at higher level of food chain.

CALICULATE CALORIC BASE LINE

CURRENT WEIGHT \times HEIGHT= CALORIC BASE LINEADD 400 CALORIES TO MAINTAIN CALORIES

CALORIC BASE LINE+400=YOUR MAINTAIN CALORIES

SUBTRACT 600 CALORIES FROM YOUR MAINTAIN CALORIES FIND THE DAILY CALORIES IN TAKE

6.many metals like silver,copper,gold,aluminium are the good conductors of heat.

materials that make good insulators do not have the ability to conduct electricity,such as glass

materials made of many different atoms are good insulators.atoms bond to gather not only with same atoma also with different atoms

7.in a hot body the molecules are in a stete of kinetic motion

and in cold body the molecules moves slowly

when hot and cold bodies in thermal contact the fast moving molecules transfer heat energy to the slowly moving molecules

entropy increases free energy decreases